# Towards Domain Adaptive Neural Contextual Bandits

### v2

@@ A.1 Proofs for Lemmas
$\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\leq\epsilon_{\mathcal{S}}(h)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)+\sum^{N}_{i=1}\alpha\cdot\kappa_{i}^{\mathcal{S}}\right),$
@@
$\equiv$ $\epsilon_{\mathcal{S}}(h)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)+\sum^{N}_{i=1}\alpha\cdot\kappa_{i}^{\mathcal{S}}\right),$
@@ A.2 Proof for the Target Regret Bound
$\quad+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)}_{Predicted~{}Rewards}},$ (16)
@@
$+\mathit{R}_{S}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\epsilon_{\mathcal{S}}(h)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)+\sum^{N}_{i=1}\alpha\cdot\kappa_{i}^{\mathcal{S}}\right)}$
@@ Appendix D Implementation Details
Appendix E Total Ablation Studies
@@ Appendix F Limitation
Intuitive Perspective. The bandit algorithm indeed enjoys interpretability and the accuracy of its closed-form updates, but this is true only because it typically employs a linear model.
@@ Appendix G Discussions
G.1 Difference between Bandit and Classification Settings
@@
G.2 Importance of Bandits and DABand
@@ G.3 Novelties Restatement
G.4 Discussion on Zero-shot Target Regret Bound
G.4.1 Significance of Theorem 3.1
@@ G.4 Discussion on Zero-shot Target Regret Bound
  • Predicted Reward. $\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\widehat{\theta},\widehat{\phi}(x_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}$, which serves as a regularization term to regularize the model to avoid overestimating rewards.
The results of the ablation study in Table 4 in Sec. 5 highlight the significance of these two terms. Moreover, we would also like to highlight the key novelty/challenges in deriving Theorem 3.1:
  • Generalization of $\mathcal{H}\Delta\mathcal{H}$ Distance to Regression. In the original multi-domain generalization bound [1], the $\mathcal{H}\Delta\mathcal{H}$ divergence between source and target domains is derived for classification models. However, contextual bandit is essentially a reward regression problem, and therefore necessitates generalizing the $\mathcal{H}\Delta\mathcal{H}$ distance from the classification case to the regression case. This presents a significant technical challenge, and leads to a series of modification in the proof.
  • Decomposing the Target Regret into the Source Regret with Other Terms. The subsequent challenge involves decomposing the regret bound for the target domain (i.e., the target regret) into a source regret term and other terms to upper-bound the target regret. This is necessary because we do not have access to feedback/reward from the target domain (we only have access to feedback/reward in the source domain). This process requires a nuanced understanding of the interplay between source and target domain dynamics within our model’s framework.
G.4.2 Testing Phase rather than Traditional Bandit Learning Settings
@@
G.4.3 Scale of Source Regret and Predicted Rewards
Similar to Neural-LinUCB, in our DABand, the true reward is restricted to the range of $[0,1]$ (as we mentioned in Sec. 3.1); therefore it will not make the source regret unbounded.
Furthermore, our DABand algorithm tries to minimize the source regret while aligning the source and target domains. The minimization of the source regret is theoretically guaranteed, as discussed in recent work such as Neural-LinUCB (Xu et al., 2020).
G.5 Are the Comparisons Fair?
@@
G.6 Why Minimize the Predicted Reward on the Source Domain?
@@
G.7 Clarification of DABand’s Contributions
@@
G.9 How the Terms in the Regret Decay as N increases.
@@
G.10 Understanding the Bound if the Source and Target Domains are equivalent
@@
G.11 Clarification on the Problem Setting.

### v3

@@ A.1 Proofs for Lemmas
$\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a_{i}^{*}}^{\mathcal{S}})\rangle-\langle\theta^{*},\phi^{*}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\leq\epsilon_{\mathcal{S}}(h)+\sum^{N}_{i=1}\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)+\alpha\cdot\kappa_{i}^{\mathcal{S}}\right),$
@@
$\equiv$ $\epsilon_{\mathcal{S}}(h)+\sum^{N}_{i=1}\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|+\alpha\cdot\kappa_{i}^{\mathcal{S}}\right),$
@@ A.2 Proof for the Target Regret Bound
$\quad+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\sum^{N}_{i=1}\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)}_{Predicted~{}Rewards}},$
@@
$+\mathit{R}_{S}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\epsilon_{\mathcal{S}}(h)+\sum^{N}_{i=1}\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|+\alpha\cdot\kappa_{i}^{\mathcal{S}}\right)}$
@@ Appendix D Implementation Details
Appendix E Full Ablation Studies
@@ Appendix F Limitation
Intuitive Perspective. The bandit algorithm indeed enjoys interpretability (Wang et al., 2024b; a) and the accuracy of its closed-form updates (Li et al., 2010), but this is true only because it typically employs a linear model.
@@ Appendix G Discussions
G.1 Sublinearity for the Target Regret Bound
Figure 2: Sub-linearity for each term in Eqn. (5).
To see the data divergence term is sub-linear, note that our target regret bound in Eqn. (5) can be divided into two parts:
  • Total Source Regret: In Neural-LinUCB (NLinUCB) (Xu et al., 2020), this is shown to be $O(\sqrt{N})$, which is sub-linear.
  • Sum of Other Terms (Including the Data Divergence): These terms can be directly optimized to convergence and minimized to a small value using SGD variants such as Adam. As shown in Corollary 4.2 of the Adam paper (Diederik, 2014), the Adam optimizer enjoys an average regret $R(N)/N$ of $O(\tfrac{1}{\sqrt{N}})$ (here $N$ is equivalent to $T$ in the Adam paper), which leads to a sub-linear total regret $O(\sqrt{N})$. Since all other terms are directly optimized using Adam, they also enjoy a sub-linear regret.
Therefore, our target regret bound is also sub-linear. Moreover:
  • Key Difference between the Source Regret and Other Terms. Note that in our target regret bound in Eqn. (5), the source regret term $\mathit{R}_{\mathcal{S}}$ increases monotonically with respect to $N$ and cannot be directly minimized. In contrast, all other terms, e.g., the data divergence term $N\cdot\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})$ can be directly minimized since all related data is training data and is already known. This is why we minimize the corresponding loss terms like Eqn. (7),  Eqn. (8) and Eqn. (9) during training.
  • The data divergence term can be minimized to a small value as $N\to\infty$. For example, if the source and target domains are perfectly aligned after our training, the data divergence becomes 00. Therefore, the cumulative sum of the data divergence term over all rounds will be sub-linear.
G.2 Difference between Bandit and Classification Settings
@@
G.3 Importance of Bandits and DABand
@@ G.4 Novelties Restatement
G.5 Discussion on Zero-shot Target Regret Bound
G.5.1 Significance of Theorem 3.1
@@ G.5 Discussion on Zero-shot Target Regret Bound
  • Predicted Reward. $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\sum^{N}_{i=1}\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)$, which serves as a regularization term to regularize the model to avoid overestimating rewards.
The results of the ablation study in Table 4 in Sec. 5 highlight the significance of these two terms. Please also refer to “Technical Novelty” of Sec. G.4 above for discussion on key novelty/challenges in deriving Theorem 3.1.
G.5.2 Testing Phase rather than Traditional Bandit Learning Settings
@@
G.5.3 Scale of Source Regret and Predicted Rewards
Similar to Neural-LinUCB (NLinUCB), in our DABand, the true reward is restricted to the range of $[0,1]$ (as we mentioned in Sec. 3.1); therefore it will not make the source regret unbounded.
Furthermore, our DABand algorithm tries to minimize the source regret while aligning the source and target domains. The minimization of the source regret is theoretically guaranteed, as discussed in recent work such as Neural-LinUCB (NLinUCB) (Xu et al., 2020).
G.6 Are the Comparisons Fair?
@@
G.7 Why Minimize the Predicted Reward on the Source Domain?
@@
G.8 Clarification of DABand’s Contributions
@@
G.10 How the Terms in the Regret Decay as N increases.
@@
G.11 Understanding the Bound if the Source and Target Domains are equivalent
@@
G.12 Clarification on the Problem Setting.